### v2

# Forecasting the U.S. Renewable-Energy Mix with an ALR-BDARMA Compositional Time-Series Framework

Abstract
Accurate forecasts of the U.S. renewable‑generation mix are essential for planning transmission upgrades, sizing storage, and setting balancing‑market rules. We introduce a Bayesian Dirichlet ARMA model (BDARMA) tailored to monthly shares of hydro, geothermal, solar, wind, wood, municipal waste, and biofuels from January 2010 through January 2025. The mean vector is modeled with a parsimonious VAR(2) in additive‑log‑ratio space, while the Dirichlet concentration parameter follows an intercept plus ten Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion.
Forecast performance is assessed with a 61‑split rolling‑origin experiment that issues twelve‑month density forecasts from January 2019 to January 2024. Compared with three alternatives—a Gaussian VAR(2) fitted in transform space, a seasonal naïve that repeats last year’s proportions, and a drift‑free ALR random walk—BDARMA lowers the mean continuous‑ranked probability score by 15–60 %, achieves component‑wise 90 % interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward. These results highlight BDARMA’s ability to deliver sharp, well‑calibrated probabilistic forecasts for multivariate renewable‑energy shares without sacrificing point precision.
Keywords: compositional time series; Dirichlet state‑space; Bayesian forecasting; renewable‑energy mix; seasonality.
1 Introduction
Electric‑sector decarbonisation hinges not only on expanding renewable output but also on anticipating how the mix of generation technologies will evolve. Hydropower, wind, solar, biomass and geothermal differ sharply in marginal cost, intermittency and siting constraints; reliable medium‑term mix forecasts therefore shape transmission expansion, storage sizing and market design (International Energy Agency, 2024; U.S. Energy Information Administration, 2024; U.S. Energy Information Administration, 2025). Renewables already supply about one‑fifth of U.S. utility‑scale electricity and their share is expected to double before 2050, making the coherence and accuracy of share forecasts more important than ever.
Why compositional methods?
Shares are compositional: they are bounded between zero and one and must sum to unity. Forecasting each component in isolation, as is common with univariate ARIMA or machine‑learning regressions (Panapakidis and Dagoumas, 2016; Chen et al., 2017), yields incoherent predictions that may turn negative or exceed 100 % (Hyndman and Athanasopoulos, 2018). Aitchison’s log‑ratio geometry provides a principled fix (Aitchison, 1986). Early multivariate illustrations, such as the VAR for geological compositions in Billheimer et al. (2001), and the state‑space model of Snyder et al. (2017), demonstrate that standard Gaussian machinery works once data are mapped to real space. Still, Gaussian log‑ratio models often overstate predictive dispersion and ignore seasonally varying volatility.
Simplex‑native likelihoods.
Recent research therefore models the composition itself. A Dirichlet ARMA process was proposed by Zheng et al. (2017), a dynamic Dirichlet–multinomial filter by Koopman et al. (2023), and a deep hierarchical Dirichlet forecaster by Das et al. (2023). In the cross‑sectional domain, Morais et al. (2018) show that Dirichlet and compositional regression can outperform traditional attraction models when explaining brand market shares, underscoring the versatility of simplex‑based methods. A direct antecedent to the present study is the Bayesian Dirichlet ARMA (BDARMA) framework introduced by Katz et al. (2024), and subsequently explored with shrinkage priors for trading‑sector shares by Katz et al. (2025), which we adopt here as the data model for a new application to the U.S. renewable‑energy mix.
Evidence from energy studies.
Applications to energy shares remain limited but growing. Compositional VAR and ARMA models, and more recently regional optimization studies, have been used to project national and sub‑national energy structures in China, the USA and Canada (Wei et al., 2021; He et al., 2022; Xu et al., 2024; Xiao and Li, 2023). Grey‑system and hybrid approaches, such as adaptive discrete grey models and MGM–BPNN–ARIMA designs for broad‑mix or bio‑energy forecasting, further boost accuracy while respecting the simplex constraint (Qian et al., 2022; Zhang et al., 2022; Suo et al., 2024). Machine‑learning work such as the LSTM study by Ma et al. (2018) and logistic growth analysis of U.S. energy trajectories by Harris et al. (2018) underline the need to tame nonlinearities, but they still rely on ad‑hoc renormalization.
Contribution.
We apply the Bayesian Dirichlet ARMA framework to the seven‑component U.S. renewable‑energy mix measured monthly from 2010 to 2024, a data set with pronounced seasonality and secular trends hitherto unaddressed in the Dirichlet literature. Forecast skill is benchmarked against three alternatives: a Gaussian VAR(2) in additive‑log‑ratio space with identical Fourier dummies, a seasonal naïve that repeats the mix observed twelve months earlier, and a drift‑free ALR random walk. A 61‑split rolling protocol produces 732 out‑of‑sample density forecasts and shows that the Dirichlet model attains the lowest continuous‑ranked probability score at every horizon while matching the VAR’s point accuracy. A full‑sample forecast issued in January 2025 encloses a trajectory where wind and solar exceed one‑third of renewable generation by early 2026, with credible bands roughly half as wide as those implied by the VAR—information directly relevant for transmission and storage planning.
The remainder of the paper is organized as follows. Section 2 describes the EIA data and seasonal covariates. Section 3 presents the BDARMA and benchmark models. Section 4 details the rolling evaluation protocol and scoring rules. Results are discussed in Section 5, and Section 6 concludes with policy implications and avenues for future research.
2 Data
The empirical analysis relies on the EIA monthly renewable‑energy consumption data set. We retain $T=181$ consecutive months from January 2010 through January 2025. Each observation is a seven‑part composition $\mathbf{y}_{t}=(y_{t,\text{hyd}},y_{t,\text{geo}},y_{t,\text{sol}},y_{t,\text{win}},y_{t,\text{woo}},y_{t,\text{was}},y_{t,\text{bio}})^{\top}\in\mathcal{S}_{7},$ where shares are obtained by dividing each raw series by their monthly total.
Additive‑log‑ratio (ALR) coordinates.
Throughout we analyse the seven‑part composition in additive‑log‑ratio form
$\mathbf{e}_{t}=\operatorname{alr}(\mathbf{y}_{t})=\bigl{(}\log\tfrac{y_{t,\text{hyd}}}{y_{t,\text{bio}}},\log\tfrac{y_{t,\text{geo}}}{y_{t,\text{bio}}},\log\tfrac{y_{t,\text{sol}}}{y_{t,\text{bio}}},\log\tfrac{y_{t,\text{win}}}{y_{t,\text{bio}}},\log\tfrac{y_{t,\text{woo}}}{y_{t,\text{bio}}},\log\tfrac{y_{t,\text{was}}}{y_{t,\text{bio}}}\bigr{)}^{\top}\in\mathbb{R}^{6},$ (1)
where biofuels serve as the common denominator (reference part). The inverse map $\operatorname{alr}^{-1}:\mathbb{R}^{6}\!\to\!\mathcal{S}_{7}$ restores a share vector via $y_{t,j}=\exp(e_{t,j})\Bigl{[}1+\sum_{k=1}^{6}\exp(e_{t,k})\Bigr{]}^{-1}$ for $j\leq 6$ and $y_{t,\text{bio}}=\bigl{[}1+\sum_{k=1}^{6}\exp(e_{t,k})\bigr{]}^{-1}$. We write $e_{t,j}$ for the $j$-th ALR coordinate and collect them as $e_{1},\dots,e_{6}$ when no time index is needed.
2.1 Exploratory data analysis
Figure 1 highlights two macro‑patterns in the sample: (i) pronounced, asymmetric intra‑annual seasonality and (ii) a medium‑run reallocation of market share from hydro to wind and solar.
Figure 1: Monthly U.S. renewable‑energy mix, 2010–2025. Hydro loses ground, while wind and solar expand rapidly. Seasonality is most pronounced in hydro (spring runoff) and wind (winter–spring peak). Areas are stacked so each month sums to 100%.
Seasonality.
Panel (a) of Figure 2 shows component‑wise box plots of monthly shares for 2010–2024; panel (b) traces the mean intra‑year profile. Hydro exhibits the largest seasonal swing, peaking in April–May and troughed in late summer while wind follows a bimodal winter/autumn pattern and solar the mirror image with a July plateau. Biomass, geothermal and waste are comparatively flat, with median intra‑year movements below 1 pp.
Cross‑source dependence.
Figure 3 (a) plots the correlation matrix of the six ALR coordinates $e_{1},\dots,e_{6}$ (biofuels as reference). Solar and wind move almost one‑for‑one relative to biofuels ($\rho_{e_{5},e_{6}}=0.97$), whereas hydro and wind are strongly anti‑correlated ($\rho\approx-0.86$). Panel (b) confirms these pair‑wise relations are non‑linear, displaying the characteristic banana‑shaped clouds induced by log‑ratio geometry.
Distributional heterogeneity.
Table 1 shows wide dispersion differences: hydro ranges from $7.5\%$ to $20.3\%$ (SD $=2.6$ pp), geothermal is quasi‑deterministic (SD $=0.17$ pp) and wood the most volatile component (SD $=5.1$ pp).
Lag‑order diagnostics.
To determine the minimum dynamic order in ALR space we fitted VAR(1) and VAR(2) models and applied Ljung–Box and Hosking portmanteau tests to the residuals (Table 2). A Ljung–Box residual diagnostic rejects the white‑noise null for coordinate $e_{3}$ under VAR(1) (p ¡ 0.001), whereas no coordinate is rejected under VAR(2) (smallest p‑value=.14). The residual ACF panels in Figure 4 show that the prominent spikes at lags 1–2 present under VAR(1) vanish when the second lag is added. At the system level, the portmanteau statistic at horizon 12 remains marginally significant; adding centered monthly dummies reduces $\chi^{2}$ from 628 to 431 ($p=0.006$). Because VAR(2) is the smallest specification to clear all short‑run autocorrelation and further lags inflate the parameter count without material gain, we adopt a VAR(2) mean and address any residual seasonality through exogenous Fourier terms.
(a) Seasonal box plots by component
(b) Average intra‑year profile
Figure 2: Seasonal variation in renewable‑energy shares, 2010–2024. Boxes span the inter‑quartile range; black bars mark the median. Means in panel (b) highlight opposing hydro/solar and hydro/wind peaks.
(a) Correlation matrix of ALR coordinates
(b) Pairwise ALR scatter plots
Figure 3: Cross‑source dependence in additive‑log‑ratio space. Positive (red) and negative (blue) correlations in panel (a) exceed 0.9 in absolute magnitude; scatter plots in panel (b) reflect the non‑linear shape induced by the simplex geometry.
Table 1: Component means and dispersion, 2010–2024 (percent of total renewables).
Hydro Geo Solar Wind Wood Waste Bio
Mean (%) 13.0 1.64 5.20 12.2 30.3 6.51 31.2
SD (%) 2.63 0.17 3.87 4.65 5.10 1.15 1.86
Q1 (%) 10.9 1.50 1.91 8.48 26.4 5.54 30.1
Q3 (%) 14.8 1.77 7.53 15.1 35.1 7.48 32.5
Min (%) 7.51 1.28 0.73 4.18 19.7 4.05 22.0
Max (%) 20.3 2.03 16.3 22.5 39.1 8.31 35.4
Figure 4: Residual autocorrelation by ALR coordinate: red=VAR(1); blue=VAR(2). Adding the second lag removes the large spikes at lags 1–2.
Table 2: Residual diagnostic statistics (lags 1–2 for Ljung–Box, horizon 12 for Hosking portmanteau).
Model Test $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$
VAR(1) Ljung–Box $p$ 0.99 0.79 0.00 0.08 0.49 0.20
VAR(2) Ljung–Box $p$ 0.99 0.98 0.47 0.73 0.57 0.14
Portmanteau $\chi^{2}$ / $p$ 628/<0.001628\;/\;<0.001628 / < 0.001 (VAR(2))
Figure 5: ACF of squared residuals for the preferred VAR(2) + season model. Modest spikes in $e_{2}$ and $e_{5}$ motivate the component‑wise GARCH precision adopted later.
Because the data exhibit markedly different seasonal amplitudes, strong yet uneven cross‑correlations, and component‑specific dispersion and because the residual diagnostics show that two lags are the minimum needed for whiteness, we adopt a specification built on three mutually reinforcing elements: a second‑order vector autoregressive mean in ALR space to capture short‑run dynamics, a single seasonal precision curve that modulates uncertainty across the calendar year for every component, and a Dirichlet likelihood that both imposes the compositional sum‑to‑one constraint and allows each source’s variance to remain distinct.
Software environment.
All computations were carried out in R 4.3.2 (R Core Team, 2024) with Stan 2.33 via the rstan interface (Goodrich et al., 2024; Stan Development Team, 2023). Data wrangling, graphics and tables relied on tidyverse (Wickham et al., 2019), lubridate (Grolemund and Wickham, 2011), janitor (Firke, 2023), scales (Wickham and Seidel, 2019), patchwork (Pedersen, 2025), ggcorrplot (Kassambara, 2022), GGally (Di Cook et al., 2021) and kableExtra (Zhu, 2021). Compositional methods used compositions (van den Boogaart and Tolosana‐Delgado, 2024) and the transport package for Aitchison norms (Schuhmacher et al., 2020). Time‑series estimation and testing employed vars (Pfaff, 2008), FinTS (Pfaff, 2024) and MTS (Tsay, 2023).
3 Forecasting model
Let the monthly renewable‑energy mix be the $J\!=\!7$-component composition
$\mathbf{y}_{t}=\bigl{(}y_{t,\text{hyd}},y_{t,\text{geo}},y_{t,\text{sol}},y_{t,\text{win}},y_{t,\text{woo}},y_{t,\text{was}},y_{t,\text{bio}}\bigr{)}^{\!\top}\in\mathcal{S}_{7},\;t=1,\dots,T.$
Biofuels ($j^{\!*}=7$) serve as reference part in every additive‑log‑ratio (ALR) transform that follows.
Dirichlet observation layer.
We model $\mathbf{y}_{t}$ as a Dirichlet distributed whose parameter vector factorizes into a simplex‑valued mean $\boldsymbol{\mu}_{t}$ and a positive precision scalar $\varphi_{t}$:
$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\varphi_{t}\;\sim\;\operatorname{Dirichlet}\bigl{(}\varphi_{t}\,\boldsymbol{\mu}_{t}\bigr{)},\qquad\boldsymbol{\mu}_{t}\in\mathcal{S}_{7},\;\varphi_{t}>0.$ (2)
Mean vector in ALR space.
Let $\boldsymbol{\eta}_{t}=\operatorname{alr}(\boldsymbol{\mu}_{t})\in\mathbb{R}^{J-1}$; for $J=7$ this is a six‑vector $\boldsymbol{\eta}_{t}=(\eta_{t1},\dots,\eta_{t6})^{\!\top}$ of log‑ratios against biofuels. Its inverse is
$\mu_{tj}=\frac{\exp(\eta_{tj})}{1+\sum_{k=1}^{6}\exp(\eta_{tk})}\;(j\leq 6),\qquad\mu_{t,j^{\!*}}=\Bigl{[}1+\textstyle\sum_{k=1}^{6}\exp(\eta_{tk})\Bigr{]}^{-1}.$
Seasonal precision curve.
Calendar variation in forecast dispersion is captured by letting the log‑precision depend on an intercept and ten Fourier harmonics:
$\log\varphi_{t}=\mathbf{f}_{t}^{\!\top}\boldsymbol{\gamma},\quad\mathbf{f}_{t}=(1,\mathbf{g}_{t}^{\!\top})^{\!\top},\;\mathbf{g}_{t}=\bigl{(}\sin\tfrac{2\pi t}{12},\cos\tfrac{2\pi t}{12},\ldots,\sin\tfrac{10\pi t}{12},\cos\tfrac{10\pi t}{12}\bigr{)}^{\!\top},\quad\boldsymbol{\gamma}\in\mathbb{R}^{11}.$ (3)
State (mean) dynamics.
Short‑run cross‑technology interactions are modelled with a second‑order vector autoregression process in ALR space:
$\boldsymbol{\eta}_{t}=\mathbf{X}_{t}\boldsymbol{\beta}+\mathbf{A}_{1}\bigl{(}\boldsymbol{\eta}_{t-1}-\mathbf{X}_{t-1}\boldsymbol{\beta}\bigr{)}+\mathbf{A}_{2}\bigl{(}\boldsymbol{\eta}_{t-2}-\mathbf{X}_{t-2}\boldsymbol{\beta}\bigr{)},\qquad\mathbf{X}_{t}=I_{J-1}\otimes\mathbf{f}_{t}^{\!\top},$ (4)
where (i) $\mathbf{A}_{1},\mathbf{A}_{2}\in\mathbb{R}^{6\times 6}$ are AR coefficient matrices; (ii) $\mathbf{X}_{t}$ block‑replicates the $11$-vector $\mathbf{f}_{t}$ across the six ALR coordinates, giving $\mathbf{X}_{t}\in\mathbb{R}^{6\times 66}$; and (iii) $\boldsymbol{\beta}\in\mathbb{R}^{66}$ contains component‑specific regression slopes for the seasonal dummies.
Prior specification and inference.
Each scalar element of $\mathbf{A}_{1},\mathbf{A}_{2},\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ receives an independent $\mathcal{N}(0,1)$ prior. Posterior inference proceeds via Hamiltonian Monte Carlo (four chains; 500 warm‑up and 500 retained iterations per chain) in Stan, yielding 2 00020002\,0002 000 draws that underpin all density‑forecast evaluations presented later.
3.1 Transform‑space VAR(2) (tVAR(2))
Working in ALR coordinates,
$\boldsymbol{\eta}_{t}=\mathbf{F}_{1}\boldsymbol{\eta}_{t-1}+\mathbf{F}_{2}\boldsymbol{\eta}_{t-2}+\mathbf{X}_{t}\boldsymbol{\delta}+\boldsymbol{\varepsilon}_{t},\qquad\boldsymbol{\varepsilon}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}).$
Parameters $(\mathbf{F}_{1},\mathbf{F}_{2},\boldsymbol{\delta},\boldsymbol{\Sigma})$ are estimated by ordinary least squares with the same seasonal regressors $\mathbf{X}_{t}$. Multi‑step forecasts are generated under the Gaussian innovation assumption and mapped back with ${\textnormal{alr}}^{-1}$.
3.2 Additive‑log‑ratio random walk (ALR–RW)
A drift‑free benchmark sets each future ALR vector equal to the most recent observation: $\boldsymbol{\eta}_{t+h\mid t}=\boldsymbol{\eta}_{t}.$ Back‑transformation yields a single point forecast with zero predictive spread.
3.3 Seasonal naïve copy‑last‑year (S‑NAIVE)
The seasonal naïve copies the composition observed 12 months earlier: $\mathbf{y}_{t+h\mid t}=\mathbf{y}_{t+h-12}.$
These four specifications exploit the same information set but differ in how they propagate seasonality, cross‑technology dependence and uncertainty. Section 4 details the rolling protocol used to compare their point and density‑forecast performance.
4 Forecast–evaluation protocol
[…]
4.1 Generating predictive distributions
All four competitors are evaluated on Monte‑Carlo samples of equal size $M=2{,}000$ so that scoring rules are comparable.
BDARMA.
For every origin $s$ we retain the $M$ posterior draws $\{\boldsymbol{\theta}^{(m)}\}_{m=1}^{M}$ returned by the Hamiltonian Monte‑Carlo sampler. Each draw is propagated through the deterministic state equation (4) for $h=1{:}H$ steps, producing the latent mean $\boldsymbol{\mu}^{(m)}_{s,h}$; a single realisation $\mathbf{y}^{(m)}_{s,h}\sim\text{Dirichlet}(\phi^{(m)}_{s,h}\boldsymbol{\mu}^{(m)}_{s,h})$ is then generated from the observation density (LABEL:eq:obs). The empirical set $\mathcal{P}^{\text{BDARMA}}_{s,h}=\{\mathbf{y}^{(m)}_{s,h}\}_{m=1}^{M}$ constitutes the predictive distribution.
tVAR(2).
Let $\widehat{\boldsymbol{\eta}}_{s,h}$ and $\widehat{\mathbf{V}}_{s,h}$ be, respectively, the conditional mean and covariance of the Gaussian forecast for the ALR vector at horizon $h$. We draw $\boldsymbol{\eta}^{(m)}_{s,h}\sim\mathcal{N}(\widehat{\boldsymbol{\eta}}_{s,h},\widehat{\mathbf{V}}_{s,h})$, transform with $\text{alr}^{-1}$, and obtain $\mathcal{P}^{\text{tVAR}}_{s,h}=\{\text{alr}^{-1}(\boldsymbol{\eta}^{(m)}_{s,h})\}_{m=1}^{M}.$ Because the back‑transformation preserves the unit sum.
ALR random walk (ALR–RW).
The point forecast is the last observed ALR vector $\boldsymbol{\eta}_{s,0}$. To give the model a distribution that can be scored with CRPS we set $\boldsymbol{\eta}^{(m)}_{s,h}=\boldsymbol{\eta}_{s,0}$ for every $m$ and define $\mathcal{P}^{\text{RW}}_{s,h}=\{\text{alr}^{-1}(\boldsymbol{\eta}_{s,0})\}_{m=1}^{M}.$ The resulting cloud is degenerate but has the same cardinality $M$.
Seasonal naïve (S‑NAIVE).
For each horizon $h$ we copy the composition observed exactly one year earlier, $\mathbf{y}_{s-12+h}$. As with the random walk we replicate this deterministic vector $M$ times, $\mathcal{P}^{\text{S–NAIVE}}_{s,h}=\{\mathbf{y}_{s-12+h}\}_{m=1}^{M}.$
4.2 Scoring rules
Denote by $\mathbf{y}_{s,h}$ the realized share vector at lead $h$ originating from window $s$. Two proper scoring rules are applied.
Continuous‑ranked probability score.
Writing $\|\mathbf{a}\|_{1}=\sum_{j=1}^{7}|a_{j}|$ for the $\ell_{1}$ norm, the sample‑based estimator is
$\operatorname{CRPS}_{s,h}(\mathcal{P})=\frac{1}{M}\sum_{m=1}^{M}\bigl{\|}\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{s,h}\bigr{\|}_{1}-\frac{1}{2M^{2}}\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\bigl{\|}\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{\,s,h}^{(m^{\prime})}\bigr{\|}_{1},$ (5)
where $\mathcal{P}$ denotes the empirical draw set for the model in question. The CRPS integrates the Brier score over all threshold events inside the simplex and is measured in share points.
Aitchison root‑mean‑square error.
Let $\widehat{\boldsymbol{\mu}}_{s,h}=M^{-1}\sum_{m}\mathbf{y}_{\,s,h}^{(m)}$ be the posterior mean. With the centred log‑ratio $\operatorname{clr}(\mathbf{p})=(\log p_{1}/g,\ldots,\log p_{7}/g)$ and geometric mean $g=(\prod_{j=1}^{7}p_{j})^{1/7}$, the point‑forecast error is
$\operatorname{RMSE}_{s,h}=\bigl{\|}\operatorname{clr}(\mathbf{y}_{s,h})-\operatorname{clr}(\widehat{\boldsymbol{\mu}}_{s,h})\bigr{\|}_{2}/\sqrt{7}.$ (6)
Both (5) and (6) reduce to zero for a perfect forecast.
Interval diagnostics.
For BDARMA the 5‑th and 95‑th sample quantiles define a 90 % credible interval for each component. Coverage is tallied over all $(s,h)$ pairs.
4.3 Fixed‑origin projection
After the rolling study a single fixed‑origin forecast is produced from the complete estimation window 2010‑01 – 2025‑01 ($\tau^{\ast}=T$). Future Fourier regressors $\mathbf{f}_{T+h}$ are generated deterministically, so the only source of uncertainty is the posterior distribution of model parameters and, for tVAR(2), the Gaussian state noise.
5 Results
5.1 Forecast accuracy across horizons
Table 3 reports the mean continuous‑ranked probability score (CRPS) and Table 4 the mean Aitchison root‑mean‑square error (RMSE) for every horizon; the same information is visualised in Figures 6 and 7. When scores are aggregated over all sixty‑one rolling origins, the Bayesian‑Dirichlet ARMA (BDARMA) dominates the field. Its mean CRPS at one step ahead is  0.00450.0045\,0.00450.0045, which is roughly one quarter lower than the transform‑space VAR(2) (tVAR(2)) and more than fifty percent lower than either naïve rule. The advantage widens as the horizon lengthens; by month 12, BDARMA still improves on tVAR(2) by around twenty percent and on the seasonal naïve (S‑NAIVE) by a factor of 1.4, while the additive‑log‑ratio random walk (ALR‑RW) remains the weakest performer throughout.
Point‑forecast results tell a complementary story. Up to eight months ahead, BDARMA and tVAR(2) deliver virtually identical RMSEs (Figure 7); thereafter the Gaussian VAR displays a slight numerical edge that peaks at just one‑hundredth of an Aitchison unit. That edge, however, is achieved at the cost of substantially wider predictive intervals, as reflected in the persistently higher CRPS. Both naïve rules fare considerably worse: S‑NAIVE holds a constant RMSE of about $0.15$ and a CRPS that is almost double BDARMA’s, whereas ALR‑RW deteriorates sharply until mid‑year before converging towards the other benchmarks only because its naive dispersion collapses.
Taken together, these findings underline three practical lessons. First, seasonality matters: the harmonic regressors embedded in $\mathbf{f}_{t}$ provide an immediate and durable boost over models that ignore calendar structure. Second, coherence alone is not enough; although tVAR(2) respects the simplex, its Gaussian scale inflates forecast spread and leaves roughly one‑fifth of the potential CRPS improvement on the table. Finally, the Dirichlet precision curve in BDARMA succeeds in producing densities that are simultaneously sharp, well calibrated and competitive in point terms even at the longest horizon considered.
Figure 6: Mean CRPS by horizon, averaged over 61 rolling origins. Lower values indicate sharper and better‑calibrated densities.
Table 3: Mean CRPS across rolling origins. Boldface marks the best score for each horizon.
Horizon BDARMA tVAR(2) S‑NAIVE ALR‑RW
1 0.00449 0.00615 0.0114 0.0086
2 0.00535 0.00740 0.0114 0.0130
3 0.00582 0.00829 0.0115 0.0165
4 0.00617 0.00875 0.0116 0.0183
5 0.00650 0.00928 0.0118 0.0196
6 0.00684 0.00968 0.0119 0.0204
7 0.00713 0.0100 0.0119 0.0201
8 0.00734 0.0102 0.0119 0.0194
9 0.00756 0.0102 0.0119 0.0178
10 0.00783 0.0103 0.0118 0.0153
11 0.00817 0.0105 0.0119 0.0129
12 0.00841 0.0106 0.0120 0.0120
Figure 7: Mean Aitchison RMSE by horizon; lower values indicate more accurate point forecasts.
Table 4: Mean Aitchison RMSE across rolling origins. Boldface marks the lowest error at each horizon.
Horizon BDARMA tVAR(2) S‑NAIVE ALR‑RW
1 0.0797 0.0821 0.145 0.114
2 0.0990 0.102 0.145 0.179
3 0.111 0.114 0.145 0.228
4 0.119 0.120 0.146 0.258
5 0.125 0.126 0.148 0.274
6 0.130 0.130 0.149 0.281
7 0.135 0.134 0.148 0.279
8 0.140 0.138 0.148 0.266
9 0.145 0.141 0.148 0.240
10 0.150 0.141 0.147 0.202
11 0.157 0.142 0.147 0.164
12 0.162 0.141 0.148 0.148
5.2 Coverage of BDARMA predictive intervals
The Monte‑Carlo intervals are well calibrated. Table 5 records the proportion of split‑horizon combinations where the realised share vector falls inside the component‑wise 90% band. Coverage rises from 86% at one month to 99% at a full year; the geometric ceiling for joint inclusion is 48% at $h=1$, so the upward trend is expected. Component averages range from 86% for biofuels to 100% for geothermal (Table 6), giving an overall mean of 94.8%.
Table 5: Empirical 90 % coverage of BDARMA component intervals.
$h$ 1 2 3 4 5 6 7 8 9 10 11 12
Cov .863 .891 .907 .933 .950 .957 .963 .975 .984 .980 .983 .985
Table 6: Component‑wise BDARMA coverage ($61\times 12$ forecasts).
Hydro Geo Solar Wind Wood Waste Bio
.939 1.000 .993 .886 .954 .998 .862
5.3 Fixed‑origin comparison
Figure 8: Twelve‑month forecasts issued 2025‑01‑01 after refitting all models to the 2010–2024 sample. Blue shading denotes the BDARMA 90 % predictive interval; colored lines are posterior or plug‑in medians. Axes are free by facet.
Figure 8 shows the one‑year‐ahead trajectories for each technology by model. The central messages are straightforward.
First, BDARMA delivers not only a median path but also a well calibrated 90 % interval. For planners this means that the same forecast set supports both point estimates—needed, for instance, in generation‑mix optimization—and a quantification of downside or upside risk that is essential when determining reserve margins, storage dispatch protocols or hedging requirements. The competing methods provide no such density information, forcing practitioners to bolt on ad‑hoc error bands or run scenario analyses that may or may not respect the unit‑sum constraint.
Second, because BDARMA’s concentration parameter is allowed to vary seasonally, the width of its interval changes appreciably across components. High‑growth and strongly seasonal sources such as wind and solar retain non‑negligible uncertainty, implying that transmission and balancing services must still accommodate significant swings even as their average shares rise. In contrast, geothermal and waste show near‑deterministic bands, suggesting that policy targets focused on these technologies can be stated with greater confidence and narrower tolerances.
Third, where the medians diverge the direction is informative in its own right. The harmonic naïve inherits last year’s level and thereby under‑projects solar in the presence of an upward trend; the random walk does the opposite by ignoring seasonality entirely; and the Gaussian VAR, although often close to BDARMA in the center, cannot reveal whether any gap that remains is material relative to forecast dispersion. From a policy standpoint this matters when setting technology‑specific incentives. Under‑estimating solar growth, for example, risks under‑sizing flexible resources needed to absorb midday surpluses, whereas over‑estimating hydro stability can mask the need for additional spring reserves.
6 Conclusion
This study develops and tests a Bayesian Dirichlet ARMA model whose additive‑log‑ratio mean follows a parsimonious VAR(2) while the Dirichlet concentration parameter evolves with an intercept plus ten monthly harmonics. In a 61‑split rolling evaluation the Dirichlet‑based approach delivers the sharpest and best‑calibrated twelve‑month density forecasts for the U.S. renewable‑energy mix. Relative to a Gaussian VAR(2) with identical seasonal dummies, mean CRPS falls by roughly one‑quarter at short leads and by one‑fifth at a full year; gains over both a seasonal naïve and an additive‑log‑ratio random walk are still larger. Point accuracy measured by Aitchison RMSE is on par with the VAR through eight months and remains within 0.02 units thereafter, underscoring that the improvement is driven by better uncertainty quantification rather than a sacrifice in central tendency.
The single‑origin forecast issued for calendar year 2025 illustrates practical implications. BDARMA produces calibrated 90% intervals that tighten for low‑volatility sources such as geothermal and waste while remaining wide enough for wind and solar to capture intra‑year swings. Because the competing models offer medians only, planners relying on them must add ad‑hoc bands or scenario grids that may violate simplex coherence. In contrast, the Dirichlet framework provides a one‑stop forecast that can feed directly into capacity‑expansion studies, transmission planning, and reliability assessments, all of which depend critically on both the mean mix and its dispersion.
Methodologically the model is straightforward to extend. Exogenous drivers like fuel prices, policy shocks, regional weather indices—can be added as additional columns in the mean or precision design matrices, and the hierarchical structure naturally scales to state or utility sub‑regions. Future work may also explore coupling the Dirichlet layer with non‑linear latent dynamics or integrating the forecast densities into stochastic power‑system optimization. For the present, the results demonstrate that a modest amount of Bayesian structure, applied in the right geometric space, is enough to yield substantial predictive gains and actionable insights for the energy‑transition debate.
Acknowledgments
Sean Wilson for insightful discussions and for his invaluable assistance in developing the original BDARMA Stan code.
Code Availability
All R scripts and Stan model files used in this study are publicly available at
https://github.com/harrisonekatz/energy-compositions. All results and figures in this manuscript can be reproduced by running the scripts found in that repository.
Conflict of Interest
The authors declare no conflicts of interest and that all work and opinions are their own and that the work is not sponsored or endorsed by Airbnb

### v3

@@
 Accurate forecasts of the U.S.
-renewable‑generation mix are essential for planning transmission upgrades, sizing storage, and setting balancing‑market rules.
+renewable-generation mix are essential for planning transmission upgrades, sizing storage, and setting balancing-market rules.
 We introduce a Bayesian Dirichlet ARMA model (BDARMA) tailored to monthly shares of hydro, geothermal, solar, wind, wood, municipal waste, and biofuels from January 2010 through January 2025.
-The mean vector is modeled with a parsimonious VAR(2) in additive‑log‑ratio space, while the Dirichlet concentration parameter follows an intercept plus ten Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion.
-Forecast performance is assessed with a 61‑split rolling‑origin experiment that issues twelve‑month density forecasts from January 2019 to January 2024.
-Compared with three alternatives—a Gaussian VAR(2) fitted in transform space, a seasonal naïve that repeats last year’s proportions, and a drift‑free ALR random walk—BDARMA lowers the mean continuous‑ranked probability score by 15–60 %, achieves component‑wise 90 % interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward.
-These results highlight BDARMA’s ability to deliver sharp, well‑calibrated probabilistic forecasts for multivariate renewable‑energy shares without sacrificing point precision.
-Keywords: compositional time series; Dirichlet state‑space; Bayesian forecasting; renewable‑energy mix; seasonality.
+The mean vector is modeled with a parsimonious VAR(2) in additive-log-ratio space, while the Dirichlet concentration parameter follows an intercept plus ten Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion.
+Forecast performance is assessed with a 61-split rolling-origin experiment that issues twelve-month density forecasts from January 2019 to January 2024.
+Compared with three alternatives—a Gaussian VAR(2) fitted in transform space, a seasonal naïve that repeats last year’s proportions, and a drift-free ALR random walk—BDARMA lowers the mean continuous-ranked probability score by 15–60 %, achieves component-wise 90 % interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward.
+These results highlight BDARMA’s ability to deliver sharp, well-calibrated probabilistic forecasts for multivariate renewable-energy shares without sacrificing point precision.
+Keywords: compositional time series; Dirichlet state-space; Bayesian forecasting; renewable-energy mix; seasonality.
 1 Introduction
-Electric‑sector decarbonisation hinges not only on expanding renewable output but also on anticipating how the mix of generation technologies will evolve.
+Electric‑sector decarbonization hinges not only on expanding renewable output but also on anticipating how the mix of generation technologies will evolve.
 Hydropower, wind, solar, biomass and geothermal differ sharply in marginal cost, intermittency and siting constraints; reliable medium‑term mix forecasts therefore shape transmission expansion, storage sizing and market design (International Energy Agency, 2024; U.S.
@@
 Distributional heterogeneity.
-Table 1 shows wide dispersion differences: hydro ranges from $7.5\%$ to $20.3\%$ (SD $=2.6$ pp), geothermal is quasi‑deterministic (SD $=0.17$ pp) and wood the most volatile component (SD $=5.1$ pp).
+Table 1 shows wide dispersion differences: hydro ranges from $7.5\%$ to $20.3\%$ (SD =2.6=2.6 pp), geothermal is quasi‑deterministic (SD =0.17=0.17 pp) and wood the most volatile component (SD =5.1=5.1 $p$).
 Lag‑order diagnostics.
 To determine the minimum dynamic order in ALR space we fitted VAR(1) and VAR(2) models and applied Ljung–Box and Hosking portmanteau tests to the residuals (Table 2).
-A Ljung–Box residual diagnostic rejects the white‑noise null for coordinate $e_{3}$ under VAR(1) (p ¡ 0.001), whereas no coordinate is rejected under VAR(2) (smallest p‑value=.14).
+A Ljung–Box residual diagnostic rejects the white‑noise null for coordinate $e_{3}$ under VAR(1) (p < 0.001), whereas no coordinate is rejected under VAR(2) (smallest p‑value=.14).
 The residual ACF panels in Figure 4 show that the prominent spikes at lags 1–2 present under VAR(1) vanish when the second lag is added.
@@
 Model Test $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$
-VAR(1) Ljung–Box $p$ 0.99 0.79 0.00 0.08 0.49 0.20
-VAR(2) Ljung–Box $p$ 0.99 0.98 0.47 0.73 0.57 0.14
-Portmanteau $\chi^{2}$ / $p$ 628/<0.001628\;/\;<0.001628 / < 0.001 (VAR(2))
+VAR(1) Ljung–Box pp 0.99 0.79 0.00 0.08 0.49 0.20
+VAR(2) Ljung–Box pp 0.99 0.98 0.47 0.73 0.57 0.14
+Portmanteau $\chi^{2}$ / pp 628/<0.001628\;/\;<0.001 (VAR(2))
 Figure 5: ACF of squared residuals for the preferred VAR(2) + season model. Modest spikes in $e_{2}$ and $e_{5}$ motivate the component‑wise GARCH precision adopted later.
@@
 3 Forecasting model
-Let the monthly renewable‑energy mix be the $J\!=\!7$-component composition
+Let the monthly renewable‑energy mix be the J=7J\!=\!7-component composition
 $\mathbf{y}_{t}=\bigl(y_{t,\text{hyd}},y_{t,\text{geo}},y_{t,\text{sol}},y_{t,\text{win}},y_{t,\text{woo}},y_{t,\text{was}},y_{t,\text{bio}}\bigr)^{\!\top}\in\mathcal{S}_{7},\;t=1,\dots,T.$
@@
 We model $\mathbf{y}_{t}$ as a Dirichlet distributed whose parameter vector factorizes into a simplex‑valued mean $\boldsymbol{\mu}_{t}$ and a positive precision scalar $\varphi_{t}$:
-$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\varphi_{t}\;\sim\;\operatorname{Dirichlet}\bigl{(}\varphi_{t}\,\boldsymbol{\mu}_{t}\bigr{)},\qquad\boldsymbol{\mu}_{t}\in\mathcal{S}_{7},\;\varphi_{t}>0.$ (2)
+$y_{t}\mid\mu_{t},\phi_{t}\sim\mathrm{Dirichlet}\!\big(\phi_{t}\mu_{t}\big),\quad\mu_{t}\in\mathcal{S}_{7},\;\phi_{t}>0.$ (2)
 Mean vector in ALR space.
@@
 where (i) $\mathbf{A}_{1},\mathbf{A}_{2}\in\mathbb{R}^{6\times 6}$ are AR coefficient matrices; (ii) $\mathbf{X}_{t}$ block‑replicates the $11$-vector $\mathbf{f}_{t}$ across the six ALR coordinates, giving $\mathbf{X}_{t}\in\mathbb{R}^{6\times 66}$; and (iii) $\boldsymbol{\beta}\in\mathbb{R}^{66}$ contains component‑specific regression slopes for the seasonal dummies.
+Geometric preliminaries and evaluation mapping
+Let $y_{t}\in\mathcal{S}_{7}$ denote the share vector and $e_{t}=\operatorname{alr}(y_{t})\in\mathbb{R}^{6}$ its additive log‑ratio (ALR) coordinates with biofuels as the reference part; $\operatorname{alr}^{-1}$ restores shares (see Eq.
+(1)).
+We model the mean in ALR space (Eq.
+(4)) and obtain predictive draws in share space from the Dirichlet observation layer (Eq.
+(2)).
+Forecasts are evaluated in two complementary spaces: CRPS in share space for joint sharpness and calibration (Eq.
+(5)), and clr‑based RMSE in Aitchison geometry for point accuracy (Eq.
+(6)).
+Because space choice induces different cross‑component dependencies, reporting both clarifies where improvements arise.
+Reference‑free coordinates.
+The logistic‑normal (sometimes “ALN”) family places a Gaussian law on log‑ratio coordinates; the isometric log‑ratio (ILR) transform provides orthonormal, reference‑free coordinates with full metric equivalence on the simplex (Aitchison and Shen, 1980; Egozcue et al., 2003; Aitchison, 1986).
+This means that the DARMA data models with these three link functions are equivalent, provided the same transformation is applied to the priors.
+We retain ALR for interpretability and continuity with Eq.
+(1).
 Prior specification and inference.
 Each scalar element of $\mathbf{A}_{1},\mathbf{A}_{2},\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ receives an independent $\mathcal{N}(0,1)$ prior.
-Posterior inference proceeds via Hamiltonian Monte Carlo (four chains; 500 warm‑up and 500 retained iterations per chain) in Stan, yielding 2 00020002\,0002 000 draws that underpin all density‑forecast evaluations presented later.
+Posterior inference proceeds via Hamiltonian Monte Carlo (four chains; 500 warm‑up and 500 retained iterations per chain) in Stan, yielding 2 0002\,000 draws that underpin all density‑forecast evaluations presented later.
 3.1 Transform‑space VAR(2) (tVAR(2))
@@
 Parameters $(\mathbf{F}_{1},\mathbf{F}_{2},\boldsymbol{\delta},\boldsymbol{\Sigma})$ are estimated by ordinary least squares with the same seasonal regressors $\mathbf{X}_{t}$.
-Multi‑step forecasts are generated under the Gaussian innovation assumption and mapped back with ${\textnormal{alr}}^{-1}$.
+Multi‑step forecasts are generated under the Gaussian innovation assumption and mapped back with $\operatorname{alr}^{-1}$.
 3.2 Additive‑log‑ratio random walk (ALR–RW)
@@
 4 Forecast–evaluation protocol
-[…]
+Model comparison follows an expanding–window, rolling‑origin design that mirrors the workflow used by system operators and energy planners.
+Let $\tau_{s}$, $s=1,\dots,S$, denote the final observation included in estimation window $s$ and let $H$ denote the fixed forecast horizon ($H=12$).
+The first origin is $\tau_{1}=\text{2019‑01}$ and the last origin that still admits a twelve‑step look‑ahead is $\tau_{S}=\text{2024‑01}$, so $S=61$.
+At origin $s$ the estimation set is $\{\mathbf{y}_{t}:1\leq t\leq\tau_{s}\}$ while the verification set comprises $\{\mathbf{y}_{\,\tau_{s}+h}:h=1,\dots,H\}$.
 4.1 Generating predictive distributions
-All four competitors are evaluated on Monte‑Carlo samples of equal size $M=2{,}000$ so that scoring rules are comparable.
+All four competitors are evaluated on Monte‑Carlo samples of equal size M=2, [^0] M=2{,}000 so that scoring rules are comparable.
 BDARMA.
 For every origin $s$ we retain the $M$ posterior draws $\{\boldsymbol{\theta}^{(m)}\}_{m=1}^{M}$ returned by the Hamiltonian Monte‑Carlo sampler.
-Each draw is propagated through the deterministic state equation (4) for $h=1{:}H$ steps, producing the latent mean $\boldsymbol{\mu}^{(m)}_{s,h}$; a single realisation $\mathbf{y}^{(m)}_{s,h}\sim\text{Dirichlet}(\phi^{(m)}_{s,h}\boldsymbol{\mu}^{(m)}_{s,h})$ is then generated from the observation density (LABEL:eq:obs).
+Each draw is propagated through the deterministic state equation (4) for h=1:Hh=1{:}H steps, producing the latent mean $\boldsymbol{\mu}^{(m)}_{s,h}$; a single realization $\mathbf{y}^{(m)}_{s,h}\sim\text{Dirichlet}(\phi^{(m)}_{s,h}\boldsymbol{\mu}^{(m)}_{s,h})$ is then generated from the observation density (2).
 The empirical set $\mathcal{P}^{\text{BDARMA}}_{s,h}=\{\mathbf{y}^{(m)}_{s,h}\}_{m=1}^{M}$ constitutes the predictive distribution.
@@
 When scores are aggregated over all sixty‑one rolling origins, the Bayesian‑Dirichlet ARMA (BDARMA) dominates the field.
-Its mean CRPS at one step ahead is  0.00450.0045\,0.00450.0045, which is roughly one quarter lower than the transform‑space VAR(2) (tVAR(2)) and more than fifty percent lower than either naïve rule.
+Its mean CRPS at one step ahead is  0.0045\,0.0045, which is roughly one quarter lower than the transform‑space VAR(2) (tVAR(2)) and more than fifty percent lower than either naïve rule.
 The advantage widens as the horizon lengthens; by month 12, BDARMA still improves on tVAR(2) by around twenty percent and on the seasonal naïve (S‑NAIVE) by a factor of 1.4, while the additive‑log‑ratio random walk (ALR‑RW) remains the weakest performer throughout.

### v4

@@
 Accurate forecasts of the U.S.
-renewable-generation mix are essential for planning transmission upgrades, sizing storage, and setting balancing-market rules.
+renewable energy consumption mix are essential for planning transmission upgrades, sizing storage, and setting balancing market rules.
 We introduce a Bayesian Dirichlet ARMA model (BDARMA) tailored to monthly shares of hydro, geothermal, solar, wind, wood, municipal waste, and biofuels from January 2010 through January 2025.
-The mean vector is modeled with a parsimonious VAR(2) in additive-log-ratio space, while the Dirichlet concentration parameter follows an intercept plus ten Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion.
-Forecast performance is assessed with a 61-split rolling-origin experiment that issues twelve-month density forecasts from January 2019 to January 2024.
-Compared with three alternatives—a Gaussian VAR(2) fitted in transform space, a seasonal naïve that repeats last year’s proportions, and a drift-free ALR random walk—BDARMA lowers the mean continuous-ranked probability score by 15–60 %, achieves component-wise 90 % interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward.
-These results highlight BDARMA’s ability to deliver sharp, well-calibrated probabilistic forecasts for multivariate renewable-energy shares without sacrificing point precision.
-Keywords: compositional time series; Dirichlet state-space; Bayesian forecasting; renewable-energy mix; seasonality.
+The mean vector is modeled with a parsimonious VAR(2) in additive log ratio space, while the Dirichlet concentration parameter follows an intercept plus five Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion.
+Forecast performance is assessed with a 61 split rolling origin experiment that issues twelve month density forecasts from January 2019 to January 2024.
+Compared with three alternatives, a Gaussian VAR(2) fitted in transform space, a seasonal naive that repeats last year’s proportions, and a drift free ALR random walk, BDARMA lowers the mean continuous ranked probability score by 15 to 60 percent, achieves component wise 90 percent interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward.
+These results highlight BDARMA’s ability to deliver sharp, well calibrated probabilistic forecasts for multivariate renewable energy shares without sacrificing point precision.
+Supplementary Material: An electric power only robustness analysis aligned with the EIA STEO baseline (strict vintaging) is reported in Sections S1 and S2.
+Keywords: Compositional time series; Dirichlet state-space; Bayesian forecasting; renewable energy mix; seasonality.
 1 Introduction
@@
 Hydropower, wind, solar, biomass and geothermal differ sharply in marginal cost, intermittency and siting constraints; reliable medium‑term mix forecasts therefore shape transmission expansion, storage sizing and market design (International Energy Agency, 2024; U.S.
-Energy Information Administration, 2024; U.S.
-Energy Information Administration, 2025).
+Energy Information Administration, 2024, 2025).
 Renewables already supply about one‑fifth of U.S.
 utility‑scale electricity and their share is expected to double before 2050, making the coherence and accuracy of share forecasts more important than ever.
-Why compositional methods?
 Shares are compositional: they are bounded between zero and one and must sum to unity.
@@
 Still, Gaussian log‑ratio models often overstate predictive dispersion and ignore seasonally varying volatility.
-Simplex‑native likelihoods.
 Recent research therefore models the composition itself.
@@
 (2018) show that Dirichlet and compositional regression can outperform traditional attraction models when explaining brand market shares, underscoring the versatility of simplex‑based methods.
-A direct antecedent to the present study is the Bayesian Dirichlet ARMA (BDARMA) framework introduced by Katz et al.
-(2024), and subsequently explored with shrinkage priors for trading‑sector shares by Katz et al.
+A direct antecedent to the present study is the Bayesian Dirichlet ARMA (BDARMA) framework (Katz et al., 2024; Katz and Weiss, 2025), and subsequently explored with shrinkage priors for trading‑sector shares by Katz et al.
 (2025), which we adopt here as the data model for a new application to the U.S.
 renewable‑energy mix.
-Evidence from energy studies.
 Applications to energy shares remain limited but growing.
@@
 (2018) underline the need to tame nonlinearities, but they still rely on ad‑hoc renormalization.
-Contribution.
 We apply the Bayesian Dirichlet ARMA framework to the seven‑component U.S.
@@
 Forecast skill is benchmarked against three alternatives: a Gaussian VAR(2) in additive‑log‑ratio space with identical Fourier dummies, a seasonal naïve that repeats the mix observed twelve months earlier, and a drift‑free ALR random walk.
-A 61‑split rolling protocol produces 732 out‑of‑sample density forecasts and shows that the Dirichlet model attains the lowest continuous‑ranked probability score at every horizon while matching the VAR’s point accuracy.
-A full‑sample forecast issued in January 2025 encloses a trajectory where wind and solar exceed one‑third of renewable generation by early 2026, with credible bands roughly half as wide as those implied by the VAR—information directly relevant for transmission and storage planning.
+A 61‑split rolling protocol produces 732 out‑of‑sample density forecasts and shows that the Dirichlet model attains the strongest probabilistic performance while maintaining the VAR’s point accuracy.
+A full-sample forecast to early 2026 projects wind and solar surpassing one-third of renewable generation, providing a coherent picture for transmission and storage planning.
 The remainder of the paper is organized as follows.
@@
 Results are discussed in Section 5, and Section 6 concludes with policy implications and avenues for future research.
+A complementary robustness analysis aligned with the EIA STEO definitions is presented in the Supplementary Material (Secs.
+S1–S2).
 2 Data
@@
 We write $e_{t,j}$ for the $j$-th ALR coordinate and collect them as $e_{1},\dots,e_{6}$ when no time index is needed.
+Electric‑power–only benchmarking.
+For comparisons to the EIASTEO industry baseline we also construct an electric‑power–only view (hydro, geothermal, solar = utility‑scale + small PV, wind, wood, waste) with monthly closure to the simplex.
+Implementation details and the strict vintaging rule are in the Supplementary Material (Sec.
+S1).
 2.1 Exploratory data analysis
@@
 Areas are stacked so each month sums to 100%.
-Seasonality.
 Panel (a) of Figure 2 shows component‑wise box plots of monthly shares for 2010–2024; panel (b) traces the mean intra‑year profile.
@@
 Biomass, geothermal and waste are comparatively flat, with median intra‑year movements below 1 pp.
-Cross‑source dependence.
 Figure 3 (a) plots the correlation matrix of the six ALR coordinates $e_{1},\dots,e_{6}$ (biofuels as reference).
-Solar and wind move almost one‑for‑one relative to biofuels ($\rho_{e_{5},e_{6}}=0.97$), whereas hydro and wind are strongly anti‑correlated ($\rho\approx-0.86$).
+Solar and wind move almost one‑for‑one relative to biofuels ($\rho_{e_{3},e_{4}}=0.97$), whereas hydro and wind are strongly anti‑correlated ($\rho_{e_{1},e_{4}}\approx-0.86$).
 Panel (b) confirms these pair‑wise relations are non‑linear, displaying the characteristic banana‑shaped clouds induced by log‑ratio geometry.
-Distributional heterogeneity.
 Table 1 shows wide dispersion differences: hydro ranges from $7.5\%$ to $20.3\%$ (SD =2.6=2.6 pp), geothermal is quasi‑deterministic (SD =0.17=0.17 pp) and wood the most volatile component (SD =5.1=5.1 $p$).
-Lag‑order diagnostics.
 To determine the minimum dynamic order in ALR space we fitted VAR(1) and VAR(2) models and applied Ljung–Box and Hosking portmanteau tests to the residuals (Table 2).
@@
 Because VAR(2) is the smallest specification to clear all short‑run autocorrelation and further lags inflate the parameter count without material gain, we adopt a VAR(2) mean and address any residual seasonality through exogenous Fourier terms.
-(a) Seasonal box plots by component
-(b) Average intra‑year profile
 Figure 2: Seasonal variation in renewable‑energy shares, 2010–2024.
@@
 Means in panel (b) highlight opposing hydro/solar and hydro/wind peaks.
-(a) Correlation matrix of ALR coordinates
-(b) Pairwise ALR scatter plots
 Figure 3: Cross‑source dependence in additive‑log‑ratio space.
@@
 Portmanteau $\chi^{2}$ / pp 628/<0.001628\;/\;<0.001 (VAR(2))
-Figure 5: ACF of squared residuals for the preferred VAR(2) + season model. Modest spikes in $e_{2}$ and $e_{5}$ motivate the component‑wise GARCH precision adopted later.
-Because the data exhibit markedly different seasonal amplitudes, strong yet uneven cross‑correlations, and component‑specific dispersion and because the residual diagnostics show that two lags are the minimum needed for whiteness, we adopt a specification built on three mutually reinforcing elements: a second‑order vector autoregressive mean in ALR space to capture short‑run dynamics, a single seasonal precision curve that modulates uncertainty across the calendar year for every component, and a Dirichlet likelihood that both imposes the compositional sum‑to‑one constraint and allows each source’s variance to remain distinct.
-Software environment.
-All computations were carried out in R 4.3.2 (R Core Team, 2024) with Stan 2.33 via the rstan interface (Goodrich et al., 2024; Stan Development Team, 2023).
+Figure 5: ACF of squared residuals for the preferred VAR(2) + season model.
+[…]
+All computations were carried out in R 4.3.2 with Stan 2.33 via the cmdstanr interface.
+(Goodrich et al., 2024; Stan Development Team, 2023).
 Data wrangling, graphics and tables relied on tidyverse (Wickham et al., 2019), lubridate (Grolemund and Wickham, 2011), janitor (Firke, 2023), scales (Wickham and Seidel, 2019), patchwork (Pedersen, 2025), ggcorrplot (Kassambara, 2022), GGally (Di Cook et al., 2021) and kableExtra (Zhu, 2021).
@@
 Biofuels ($j^{\!*}=7$) serve as reference part in every additive‑log‑ratio (ALR) transform that follows.
-Dirichlet observation layer.
 We model $\mathbf{y}_{t}$ as a Dirichlet distributed whose parameter vector factorizes into a simplex‑valued mean $\boldsymbol{\mu}_{t}$ and a positive precision scalar $\varphi_{t}$:
 $y_{t}\mid\mu_{t},\phi_{t}\sim\mathrm{Dirichlet}\!\big(\phi_{t}\mu_{t}\big),\quad\mu_{t}\in\mathcal{S}_{7},\;\phi_{t}>0.$ (2)
-Mean vector in ALR space.
 Let $\boldsymbol{\eta}_{t}=\operatorname{alr}(\boldsymbol{\mu}_{t})\in\mathbb{R}^{J-1}$; for $J=7$ this is a six‑vector $\boldsymbol{\eta}_{t}=(\eta_{t1},\dots,\eta_{t6})^{\!\top}$ of log‑ratios against biofuels.
@@
 $\mu_{tj}=\frac{\exp(\eta_{tj})}{1+\sum_{k=1}^{6}\exp(\eta_{tk})}\;(j\leq 6),\qquad\mu_{t,j^{\!*}}=\Bigl[1+\textstyle\sum_{k=1}^{6}\exp(\eta_{tk})\Bigr]^{-1}.$
-Seasonal precision curve.
-Calendar variation in forecast dispersion is captured by letting the log‑precision depend on an intercept and ten Fourier harmonics:
-$\log\varphi_{t}=\mathbf{f}_{t}^{\!\top}\boldsymbol{\gamma},\quad\mathbf{f}_{t}=(1,\mathbf{g}_{t}^{\!\top})^{\!\top},\;\mathbf{g}_{t}=\bigl(\sin\tfrac{2\pi t}{12},\cos\tfrac{2\pi t}{12},\ldots,\sin\tfrac{10\pi t}{12},\cos\tfrac{10\pi t}{12}\bigr)^{\!\top},\quad\boldsymbol{\gamma}\in\mathbb{R}^{11}.$ (3)
-State (mean) dynamics.
+Calendar variation in forecast dispersion is captured by letting the log‑precision depend on an intercept and five Fourier harmonics (ten sine/cosine terms):
+$\log\phi_{t}=\mathbf{f}_{t}^{\!\top}\boldsymbol{\gamma},\quad\mathbf{f}_{t}=(1,\mathbf{g}_{t}^{\!\top})^{\!\top},\;\mathbf{g}_{t}=\bigl(\sin\tfrac{2\pi t}{12},\cos\tfrac{2\pi t}{12},\ldots,\sin\tfrac{10\pi t}{12},\cos\tfrac{10\pi t}{12}\bigr)^{\!\top},\quad\boldsymbol{\gamma}\in\mathbb{R}^{11}.$ (3)
 Short‑run cross‑technology interactions are modelled with a second‑order vector autoregression process in ALR space:
@@
 where (i) $\mathbf{A}_{1},\mathbf{A}_{2}\in\mathbb{R}^{6\times 6}$ are AR coefficient matrices; (ii) $\mathbf{X}_{t}$ block‑replicates the $11$-vector $\mathbf{f}_{t}$ across the six ALR coordinates, giving $\mathbf{X}_{t}\in\mathbb{R}^{6\times 66}$; and (iii) $\boldsymbol{\beta}\in\mathbb{R}^{66}$ contains component‑specific regression slopes for the seasonal dummies.
+With a scalar precision $\phi_{t}$, the Dirichlet implies a restricted covariance: $\mathrm{Cov}(y_{i},y_{j})=-\mu_{i}\mu_{j}/(\phi_{t}+1)$ for $i\neq j$, i.e., negative off‑diagonals of fixed shape.
+Cross‑component comovement beyond the unit‑sum constraint therefore enters through the mean dynamics rather than the observation variance.
+Extensions include generalized Dirichlet or logistic‑normal layers, or component‑specific precisions $\phi_{j,t}$ with regularization; we leave these for future work.
 Geometric preliminaries and evaluation mapping
@@
 (1).
-Prior specification and inference.
 Each scalar element of $\mathbf{A}_{1},\mathbf{A}_{2},\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ receives an independent $\mathcal{N}(0,1)$ prior.
@@
 Let $\widehat{\boldsymbol{\eta}}_{s,h}$ and $\widehat{\mathbf{V}}_{s,h}$ be, respectively, the conditional mean and covariance of the Gaussian forecast for the ALR vector at horizon $h$.
-We draw $\boldsymbol{\eta}^{(m)}_{s,h}\sim\mathcal{N}(\widehat{\boldsymbol{\eta}}_{s,h},\widehat{\mathbf{V}}_{s,h})$, transform with $\text{alr}^{-1}$, and obtain $\mathcal{P}^{\text{tVAR}}_{s,h}=\{\text{alr}^{-1}(\boldsymbol{\eta}^{(m)}_{s,h})\}_{m=1}^{M}.$ Because the back‑transformation preserves the unit sum.
+We draw $\boldsymbol{\eta}^{(m)}_{s,h}\sim\mathcal{N}(\widehat{\boldsymbol{\eta}}_{s,h},\widehat{\mathbf{V}}_{s,h})$, transform with $\text{alr}^{-1}$, and obtain $\mathcal{P}^{\text{tVAR}}_{s,h}=\{\text{alr}^{-1}(\boldsymbol{\eta}^{(m)}_{s,h})\}_{m=1}^{M}.$ Multi‑step forecasts are generated under the Gaussian‑innovation assumption and then mapped back to shares with $\operatorname{alr}^{-1}$, which preserves unit‑sum coherence.
 ALR random walk (ALR–RW).
@@
 Denote by $\mathbf{y}_{s,h}$ the realized share vector at lead $h$ originating from window $s$. Two proper scoring rules are applied.
-Continuous‑ranked probability score.
-Writing $\|\mathbf{a}\|_{1}=\sum_{j=1}^{7}|a_{j}|$ for the $\ell_{1}$ norm, the sample‑based estimator is
-$\operatorname{CRPS}_{s,h}(\mathcal{P})=\frac{1}{M}\sum_{m=1}^{M}\bigl\|\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{s,h}\bigr\|_{1}-\frac{1}{2M^{2}}\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\bigl\|\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{\,s,h}^{(m^{\prime})}\bigr\|_{1},$ (5)
-where $\mathcal{P}$ denotes the empirical draw set for the model in question. The CRPS integrates the Brier score over all threshold events inside the simplex and is measured in share points.
+Energy score (multivariate CRPS).
+Writing $\|\mathbf{a}\|_{1}=\sum_{j=1}^{7}|a_{j}|$ for the $\ell_{1}$ norm, the sample‑based energy score (ES; a multivariate generalization of the CRPS) is
+$\operatorname{ES}_{s,h}(\mathcal{P})=\frac{1}{M}\sum_{m=1}^{M}\bigl\|\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{s,h}\bigr\|_{1}-\frac{1}{2M^{2}}\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\bigl\|\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{\,s,h}^{(m^{\prime})}\bigr\|_{1}.$ (5)
+We use the $\ell_{1}$ norm so that units are “share points”; ES remains a strictly proper scoring rule under common norms.
 Aitchison root‑mean‑square error.
@@
 For BDARMA the 5‑th and 95‑th sample quantiles define a 90 % credible interval for each component. Coverage is tallied over all $(s,h)$ pairs.
+4.3 External baseline and scoring in electric‑only space
+A matched evaluation against the vintaged EIASTEO baseline in the EP‑only frame uses the same rolling origins, horizons, and scoring rules; see Supplementary Sec.
+S1 for construction and Sec.
+S2 for the horizon‑by‑horizon results (Figs.
+S1–S2; Tables S1–S2).
 4.4 Fixed‑origin projection
@@
 5.1 Forecast accuracy across horizons
-Table 3 reports the mean continuous‑ranked probability score (CRPS) and Table 4 the mean Aitchison root‑mean‑square error (RMSE) for every horizon; the same information is visualised in Figures 6 and 7.
-When scores are aggregated over all sixty‑one rolling origins, the Bayesian‑Dirichlet ARMA (BDARMA) dominates the field.
-Its mean CRPS at one step ahead is  0.0045\,0.0045, which is roughly one quarter lower than the transform‑space VAR(2) (tVAR(2)) and more than fifty percent lower than either naïve rule.
-The advantage widens as the horizon lengthens; by month 12, BDARMA still improves on tVAR(2) by around twenty percent and on the seasonal naïve (S‑NAIVE) by a factor of 1.4, while the additive‑log‑ratio random walk (ALR‑RW) remains the weakest performer throughout.
-Point‑forecast results tell a complementary story.
-Up to eight months ahead, BDARMA and tVAR(2) deliver virtually identical RMSEs (Figure 7); thereafter the Gaussian VAR displays a slight numerical edge that peaks at just one‑hundredth of an Aitchison unit.
-That edge, however, is achieved at the cost of substantially wider predictive intervals, as reflected in the persistently higher CRPS.
-Both naïve rules fare considerably worse: S‑NAIVE holds a constant RMSE of about $0.15$ and a CRPS that is almost double BDARMA’s, whereas ALR‑RW deteriorates sharply until mid‑year before converging towards the other benchmarks only because its naive dispersion collapses.
-Taken together, these findings underline three practical lessons.
-First, seasonality matters: the harmonic regressors embedded in $\mathbf{f}_{t}$ provide an immediate and durable boost over models that ignore calendar structure.
-Second, coherence alone is not enough; although tVAR(2) respects the simplex, its Gaussian scale inflates forecast spread and leaves roughly one‑fifth of the potential CRPS improvement on the table.
-Finally, the Dirichlet precision curve in BDARMA succeeds in producing densities that are simultaneously sharp, well calibrated and competitive in point terms even at the longest horizon considered.
+Tables 3 and 4 report mean CRPS and mean Aitchison RMSE by horizon; Figures 6 and 7 visualize the same quantities.
+Across sixty‑one rolling origins, the Bayesian Dirichlet ARMA (BDARMA) model attains the lowest CRPS at every horizon.
+At one month it is about one quarter lower than the transform‑space VAR(2) and more than half lower than either naïve rule, and the advantage widens with lead: by twelve months BDARMA still improves on tVAR(2) by roughly one fifth and on S‑NAIVE by about forty percent, while ALR‑RW remains weakest overall.
+Point errors give a complementary view.
+Through eight months, BDARMA and tVAR(2) yield nearly identical RMSEs; from month nine onward, the Gaussian VAR gains a small edge that peaks at roughly one hundredth of an Aitchison unit.
+That edge comes with broader predictive spreads, which shows up as persistently higher CRPS for the VAR.
+Technology‑specific interpretation.
+Component patterns matter for planning.
+Wind and hydro exhibit the largest seasonal swings, so their predictive bands are wider and more seasonally structured, informing spring runoff scheduling for hydro and winter ramping reserves for wind.
+Solar’s long‑run rise with a summer plateau produces medium‑horizon gains that help quantify midday surplus risk and storage sizing.
+Geothermal and waste behave almost deterministically, supporting narrow tolerance bands for compliance or procurement.
+Wood remains comparatively volatile across horizons, arguing for conservative hedging where biomass supply or policy constraints bind.
+These qualitative statements align with Table 6, where geothermal and waste approach perfect inclusion while wind and biofuels are harder to capture due to stronger seasonality and policy or demand variability.
 Figure 6: Mean CRPS by horizon, averaged over 61 rolling origins. Lower values indicate sharper and better‑calibrated densities.
@@
 12 0.00841 0.0106 0.0120 0.0120
-Figure 7: Mean Aitchison RMSE by horizon; lower values indicate more accurate point forecasts.
+Figure 7: Mean Aitchison RMSE by horizon. Lower values indicate more accurate point forecasts.
 Table 4: Mean Aitchison RMSE across rolling origins. Boldface marks the lowest error at each horizon.
@@
 5.2 Coverage of BDARMA predictive intervals
-The Monte‑Carlo intervals are well calibrated.
-Table 5 records the proportion of split‑horizon combinations where the realised share vector falls inside the component‑wise 90% band.
-Coverage rises from 86% at one month to 99% at a full year; the geometric ceiling for joint inclusion is 48% at $h=1$, so the upward trend is expected.
-Component averages range from 86% for biofuels to 100% for geothermal (Table 6), giving an overall mean of 94.8%.
-Table 5: Empirical 90 % coverage of BDARMA component intervals.
+The Monte Carlo intervals are well calibrated.
+Componentwise 90% coverage rises from 86% at one month to 99% by a full year (Table 5).
+By technology (Table 6), geothermal and waste approach perfect inclusion; solar and hydro are very close to nominal; wind and biofuels are lower, consistent with larger seasonal amplitude and policy‑ or demand‑driven swings.
+Table 5: Empirical 90 percent coverage of BDARMA component intervals.
 $h$ 1 2 3 4 5 6 7 8 9 10 11 12
 Cov .863 .891 .907 .933 .950 .957 .963 .975 .984 .980 .983 .985
-Table 6: Component‑wise BDARMA coverage ($61\times 12$ forecasts).
+Table 6: Componentwise BDARMA coverage (61 by 12 forecasts).
 Hydro Geo Solar Wind Wood Waste Bio
@@
 5.3 Fixed‑origin comparison
-Figure 8: Twelve‑month forecasts issued 2025‑01‑01 after refitting all models to the 2010–2024 sample.
-Blue shading denotes the BDARMA 90 % predictive interval; colored lines are posterior or plug‑in medians.
+Figure 8: Twelve‑month forecasts issued 2025-01-01 after refitting all models to the 2010–2024 sample.
+Blue shading denotes the BDARMA 90% predictive interval.
+Colored lines are posterior or plug‑in medians.
 Axes are free by facet.
-Figure 8 shows the one‑year‐ahead trajectories for each technology by model. The central messages are straightforward.
-First, BDARMA delivers not only a median path but also a well calibrated 90 % interval.
-For planners this means that the same forecast set supports both point estimates—needed, for instance, in generation‑mix optimization—and a quantification of downside or upside risk that is essential when determining reserve margins, storage dispatch protocols or hedging requirements.
-The competing methods provide no such density information, forcing practitioners to bolt on ad‑hoc error bands or run scenario analyses that may or may not respect the unit‑sum constraint.
-Second, because BDARMA’s concentration parameter is allowed to vary seasonally, the width of its interval changes appreciably across components.
-High‑growth and strongly seasonal sources such as wind and solar retain non‑negligible uncertainty, implying that transmission and balancing services must still accommodate significant swings even as their average shares rise.
-In contrast, geothermal and waste show near‑deterministic bands, suggesting that policy targets focused on these technologies can be stated with greater confidence and narrower tolerances.
-Third, where the medians diverge the direction is informative in its own right.
-The harmonic naïve inherits last year’s level and thereby under‑projects solar in the presence of an upward trend; the random walk does the opposite by ignoring seasonality entirely; and the Gaussian VAR, although often close to BDARMA in the center, cannot reveal whether any gap that remains is material relative to forecast dispersion.
-From a policy standpoint this matters when setting technology‑specific incentives.
-Under‑estimating solar growth, for example, risks under‑sizing flexible resources needed to absorb midday surpluses, whereas over‑estimating hydro stability can mask the need for additional spring reserves.
+The one‑year trajectories in Figure 8 reinforce these patterns.
+BDARMA produces calibrated bands that widen where seasonality and trend are strong and tighten where series are stable, supporting a single forecast set for both point planning and risk assessment.
+The naïve rules supply medians only.
+The Gaussian VAR yields medians close to BDARMA but cannot indicate whether any remaining gap is material relative to forecast dispersion.
+If decisions depend almost entirely on point forecasts, exogenous regressors are abundant and trusted, and speed is critical, the transform‑space VAR(2) is reasonable, especially beyond nine months where it has a slight RMSE edge.
+The seasonal naïve can suffice for steady components and as a monitoring benchmark where deviations from a baseline are the main signal.
+BDARMA remains preferable when calibrated densities, compositional coherence, and seasonal uncertainty are central to the decision.
+Sequence models such as LSTMs or GRUs can be competitive with long histories and rich covariates.
+In monthly, medium‑length settings with few components, they require careful constraints to preserve the unit sum and substantial tuning to obtain calibrated densities.
+Relative to such sequence learners and to gradient‑boosted ensembles, BDARMA offers three practical advantages for this task: (i) coherence by construction (no post‑hoc renormalization), (ii) direct density forecasts in share space rather than ad‑hoc bands, and (iii) strong performance with limited history.
+The approaches are complementary: neural or boosted summaries of weather or policy can enter BDARMA as exogenous features, and neural forecasts can be used as external signals.
+Estimation uses four chains with 1000 warm‑up and draw counts; origins and chains parallelize across cores. Design matrices grow linearly with the number of components and harmonics.
+A matched electric‑power‑only comparison against the vintaged EIASTEO baseline confirms the main findings: BDARMA leads at one month, while STEO is strongest beyond two months.
+Details are in the Supplementary Material (Sections S1–S2; Figures S1–S2; Tables S1–S2).
 6 Conclusion
-This study develops and tests a Bayesian Dirichlet ARMA model whose additive‑log‑ratio mean follows a parsimonious VAR(2) while the Dirichlet concentration parameter evolves with an intercept plus ten monthly harmonics.
-In a 61‑split rolling evaluation the Dirichlet‑based approach delivers the sharpest and best‑calibrated twelve‑month density forecasts for the U.S.
-renewable‑energy mix.
-Relative to a Gaussian VAR(2) with identical seasonal dummies, mean CRPS falls by roughly one‑quarter at short leads and by one‑fifth at a full year; gains over both a seasonal naïve and an additive‑log‑ratio random walk are still larger.
-Point accuracy measured by Aitchison RMSE is on par with the VAR through eight months and remains within 0.02 units thereafter, underscoring that the improvement is driven by better uncertainty quantification rather than a sacrifice in central tendency.
-The single‑origin forecast issued for calendar year 2025 illustrates practical implications.
-BDARMA produces calibrated 90% intervals that tighten for low‑volatility sources such as geothermal and waste while remaining wide enough for wind and solar to capture intra‑year swings.
-Because the competing models offer medians only, planners relying on them must add ad‑hoc bands or scenario grids that may violate simplex coherence.
-In contrast, the Dirichlet framework provides a one‑stop forecast that can feed directly into capacity‑expansion studies, transmission planning, and reliability assessments, all of which depend critically on both the mean mix and its dispersion.
-Methodologically the model is straightforward to extend.
-Exogenous drivers like fuel prices, policy shocks, regional weather indices—can be added as additional columns in the mean or precision design matrices, and the hierarchical structure naturally scales to state or utility sub‑regions.
-Future work may also explore coupling the Dirichlet layer with non‑linear latent dynamics or integrating the forecast densities into stochastic power‑system optimization.
-For the present, the results demonstrate that a modest amount of Bayesian structure, applied in the right geometric space, is enough to yield substantial predictive gains and actionable insights for the energy‑transition debate.
+This paper develops and evaluates a Bayesian Dirichlet ARMA (BDARMA) model for monthly U.S.
+renewable‑energy shares, with a VAR(2) mean in additive‑log‑ratio space and a seasonal Dirichlet precision.
+In a 61‑split rolling evaluation, BDARMA delivers the sharpest, best‑calibrated twelve‑month density forecasts: mean CRPS is lower than a transform‑space VAR(2) at every horizon and far below naïve rules, while point accuracy (Aitchison RMSE) matches the VAR through eight months and stays within roughly two hundredths thereafter.
+The result is improved uncertainty quantification without sacrificing the central path.
+The errors have planning implications by technology.
+Hydro and wind, our most seasonal components, retain wider, seasonally patterned bands that inform spring runoff scheduling and winter ramping reserves.
+Solar’s trend plus summer plateau yields medium‑horizon gains that tighten estimates of midday surplus risk and storage needs.
+Geothermal and waste are near‑deterministic, justifying narrow tolerance bands, while wood’s broader dispersion argues for conservative procurement where fuel and policy are more volatile.
+A balanced view also clarifies when simpler models are reasonable.
+VAR(2) can be preferred when decisions hinge almost entirely on point forecasts, exogenous regressors are abundant, and computational speed is paramount; the seasonal naïve is defensible for steady systems and as a monitoring baseline.
+By contrast, BDARMA should be the default when calibrated densities, coherence across components, and explicit seasonality matter for reserves, transmission, and storage planning.
 Acknowledgments